# Finite-temperature entanglement and coherence in asymmetric bosonic Josephson junctions

### v1

@@ II.1 The model Hamiltonian
with $Z$ being the canonical partition function. The thermal average of an operator $\hat{Q}$ is then $\langle\hat{Q}\rangle=\text{Tr}(\hat{\rho}\hat{Q})$.
@@ II.2 Quantum phase model
Following Pitaevskii and Stringari [21], the classical $H_{J}$ can then be quantized by promoting the conjugate variables $\theta$ and $\hbar k$ to operators satisfying the commutation relation $[\hat{\theta},\hbar\hat{k}]=i\hbar$. In the ‘$\theta$ representation’, $\hat{k}=-i\partial/\partial\theta$ and
@@ II.3 Semiclassical approximation
[…]
@@ III.1 Spectral decomposition
Figure 1: Spectral decomposition $\rho_{ii}=\langle|c_{i}|^{2}\rangle$ of the statistical ensemble of states for $\varepsilon=0$ as a function of $i/N$, plotted for $N=50$ and $U/J=1.0$ (solid orange line), $0.0$ (dashed green line), $-0.2$ (dashed-dotted cyan line) at different temperatures.
@@
Figure 2: Spectral decomposition $\rho_{ii}=\langle|c_{i}|^{2}\rangle$ of the statistical ensemble of states for $\varepsilon/J=0.2$ as a function of $i/N$, plotted for $N=50$ and $U/J=1.0$ (solid orange line), $0.0$ (dashed green line), $-0.2$ (dashed-dotted cyan line) at different temperatures.
[…]
@@ III.2 Thermodynamic entropy
where $\rho_{n}=e^{-\beta E_{n}}/Z$. This is written in terms of the internal energy $E=\langle\hat{H}\rangle$ and the Helmholtz free energy $F=-k_{B}T\ln Z$ as $S=(E-F)/T$, that is the usual thermodynamic relation.
@@ III.3 Entanglement entropy
For a system with $N$ particles, $S_{E}$ takes values in the interval $[0,\ln(N+1)]$. It is zero if and only if $\hat{\rho}_{R}$ describes a pure state, while it is $\ln(N+1)$ when the system is in the maximally entangled state described by $\hat{\rho}_{R}=\mathbb{1}/(N+1)$.
@@ III.4 Population imbalance
which is bounded in the interval $[-N/2,\,N/2]$. In the QPM, $\hat{k}=-i\partial/\partial\theta$ and its average can be computed as discussed in Appendix A. Both these definitions require a numerical calculation. However, in the strong tunneling regime we can obtain an analytic result using the semiclassical approach discussed in Sec. II.3; from Eq. (11) we get
@@
where erf$(x)$ denotes the error function. The slope of $k(\varepsilon)$ at $\varepsilon=0$ represents the susceptibility of the particle imbalance to the on-site energy asymmetry $\varepsilon$, and can be regarded as a sort of capacitance of the bosonic Josephson junction in the linear-response regime. Deriving Eq. (21) with respect to $\varepsilon$, we get
@@ III.5 Quantum Fisher information
and takes values in the interval $[0,1]$. For pure states, e.g. the ground state when the system is at zero temperature, a sufficient (but not necessary) condition for particle entanglement is $\mathcal{I}>1/N$ [47]. In addition to being a flag for entanglement, $\mathcal{I}$ carries information about quantum fluctuations in the relative number of particles and relative phase. Since these are conjugate variables satisfying $[\hat{\theta},\hat{k}]=i$, we can use the Fisher information to estimate a lower bound for the quantum fluctuation of the phase $\Delta\theta^{2}\equiv\langle\theta^{2}\rangle-\langle\theta\rangle^{2}$ via the uncertainty principle $\Delta\theta\Delta k\geq 1/2$, which yields $\Delta\theta^{2}\geq 1/(N^{2}\mathcal{I})$. Consequently, $\Delta\theta^{2}$ becomes extremely large as $\mathcal{I}\to 0$, while it may be as small as $1/N^{2}$ for $\mathcal{I}\to 1$.
@@
[…]
@@ III.6 Coherence visibility
$=\frac{1}{Z}\sum_{n,j=0}^{N}e^{-\beta E_{n}}\sqrt{4(1+j)(1-j/N)}c_{j+1}^{(n)}c_{j}^{(n)}$ (29)
@@ Appendix A Validity of the quantum phase model and possible improvements
The number of eigenstates $\psi_{n}(\theta)$ is in principle infinite, reflecting the fact that this approach is naturally suited to describe the system in the thermodynamic limit. Indeed, with canonical quantization the classical observable $k$, bounded in the interval $[-N/2,+N/2]$, is promoted to the unbounded operator $\hat{k}=-i\partial/\partial\theta$, and the mapping between the two can be made rigorous only in the limit $N\to\infty$. However, we can still apply the QPM to describe a system with a finite number $N$ of bosons, by introducing a cutoff on the label $n$ in such a way that the total number of eigenstates (i.e. the dimension of the Hilbert space) is equal to the dimension of the Hilbert space of (II.1), which is $N+1$. If $N$ is even, taking into account the ordering of the eigenvalues in terms of Mathieu characteristic numbers, this means that $n$ must take on the integer values from 00 to $N/2$. Thermal averages may then be computed as
@@
and $Z=\sum_{n=0}^{N/2}e^{-\beta E_{n}^{(a)}}+\sum_{n=1}^{N/2}e^{-\beta E_{n}^{(b)}}$. If instead $N$ is odd, the index $n$ would take on the integer values between 00 and $\lfloor N/2\rfloor$ in the $a$-summations, and the integer values between $1$ and $\lceil N/2\rceil$ in the $b$-summations.

### v2

@@ II.1 The model Hamiltonian
[…]
@@ II.2 Quantum phase model
Following Pitaevskii and Stringari [21], the classical $H_{J}$ can then be quantized by promoting the conjugate variables $\theta$ and $\hbar k$ to operators satisfying the commutation relation $\hat{\theta},\hbar\hat{k}]=i\hbar$. In the $\theta$ representation’, $\hat{k}=-i\partial/\partial\theta$ and
@@ II.3 Semiclassical approximation
$\langle\hat{Q}\rangle=\frac{1}{Z}\int_{-\frac{N}{2}}^{\frac{N}{2}}dk\int_{-\pi}^{\pi}d\theta\,Q(k,\theta)e^{-\beta_{\text{eff}}H_{J}(k,\theta)},$ (11)
with $Z=\int_{-N/2}^{N/2}dk\int_{-\pi}^{\pi}d\theta\,e^{-\beta_{\text{eff}}H_{J}(k,\theta)}$. It is clear that this possibility is conditioned on the existence of an observable $Q(k,\theta)$ that is the classical analogue of the operator $\hat{Q}$, which is not always the case. As we will discuss in the following sections, depending on the specific observable we are considering, the range of validity of Eq. (11) may be subject to further conditions related to the number of particles $N$ and the temperature $T$.
@@ III.1 Spectral decomposition
Figure 1: Spectral decomposition $\rho_{ii}=\langle|c_{i}|^{2}\rangle$ of the statistical ensemble of states for $\varepsilon=0$ as a function of $i/N$, plotted for $N=50$ and $U/J=1.0$ (solid orange line), $0.0$ (dashed green line), 0.2-0.2 (dashed-dotted cyan line) at different temperatures.
@@
Figure 2: Spectral decomposition $\rho_{ii}=\langle|c_{i}|^{2}\rangle$ of the statistical ensemble of states for $\varepsilon/J=0.2$ as a function of $i/N$, plotted for $N=50$ and $U/J=1.0$ (solid orange line), $0.0$ (dashed green line), 0.2-0.2 (dashed-dotted cyan line) at different temperatures.
Introducing a nonzero $\varepsilon$ explicitly breaks the left-right symmetry and lifts the quasi-degeneracy of the two lowest eigenstates (Fig. 3b). In this case no quantum phase transition occurs. Given $\varepsilon$, the left-right symmetry of $|c_{i}^{(0)}|^{2}$ is broken around a specific value of $U/J$, which is negative for weak asymmetry and positive for moderate and large asymmetry. For strong attraction the ground state tends to $|0,N\rangle$ for $\varepsilon>0$ or $|N,0\rangle$ for $\varepsilon<0$, while for strong repulsion the effect of the asymmetry is overcome, and the ground state tends in any case to $|\text{FOCK}\rangle$. At finite temperature, thermal fluctuations tend to give weight to all Fock states, attenuating the symmetry-breaking effects of $\varepsilon$ (Fig. 2).
We note that at infinite temperature, the density matrix $\hat{\rho}=\mathbb{1}/(N+1)$ describes a uniform mixture of the $N+1$ Fock states, and the quantum average of an operator $\hat{Q}$ is $\langle\hat{Q}\rangle_{\beta=0}=(N+1)^{-1}\sum_{i=0}^{N}\langle i,N-i|\hat{Q}|i,N-i\rangle$. Any diagonal operator in the Fock basis will thus average to the arithmetic mean of its eigenvalues, whereas any operator with off-diagonal character will average to zero. In the semiclassical approximation, the same average is given by $\langle\hat{Q}\rangle_{\beta=0}^{\text{cl}}=(2\pi N)^{-1}\int_{-N/2}^{N/2}dk\int_{-\pi}^{\pi}d\theta\,Q(k,\theta)$. While the two averages coincide in the large $N$ limit, at finite $N$ this is not guaranteed, but depends on the nature of the operator; in fact, they do not coincide when $\hat{Q}$ is probing genuine quantum correlations. This may place an upper bound on the range of temperatures over which the semiclassical approximation can give quantitatively accurate results at finite $N$. A concrete example will be discussed in Sec. III.5.
@@ III.2 Thermodynamic entropy
where $\rho_{n}=e^{-\beta E_{n}}/Z$. This is written in terms of the internal energy $E=\langle\hat{H}\rangle$ and the Helmholtz free energy $F=-k_{B}T\ln Z$ as S=(EF)/TS=(E-F)/T, that is the usual thermodynamic relation.
@@ III.3 Entanglement entropy
For a system with $N$ particles, $S_{E}$ takes values in the interval $0,\ln(N+1)]$. It is zero if and only if $\hat{\rho}_{R}$ describes a pure state, while it is $\ln(N+1)$ when the system is in the maximally entangled state described by $\hat{\rho}_{R}=\mathbb{1}/(N+1)$.
@@ III.4 Population imbalance
which is bounded in the interval [N/2,N/2][-N/2,\,N/2]. In the QPM, $\hat{k}=-i\partial/\partial\theta$ and its average can be computed as discussed in Appendix A. Both these definitions require a numerical calculation. However, in the strong tunneling regime we can obtain an analytic result using the semiclassical approach discussed in Sec. II.3; from Eq. (11) we get
@@
where erf(x)(x) denotes the error function. The slope of $k(\varepsilon)$ at $\varepsilon=0$ represents the susceptibility of the particle imbalance to the on-site energy asymmetry $\varepsilon$, and can be regarded as a sort of capacitance of the bosonic Josephson junction in the linear-response regime. Deriving Eq. (21) with respect to $\varepsilon$, we get
@@ III.5 Quantum Fisher information
and takes values in the interval $[0,1]$. For pure states, e.g. the ground state when the system is at zero temperature, a sufficient (but not necessary) condition for particle entanglement is $\mathcal{I}>1/N$ [47]. In addition to being a flag for entanglement, $\mathcal{I}$ carries information about quantum fluctuations in the relative number of particles and relative phase. Since these are conjugate variables satisfying $\hat{\theta},\hat{k}]=i$, we can use the Fisher information to estimate a lower bound for the quantum fluctuation of the phase $\Delta\theta^{2}\equiv\langle\theta^{2}\rangle-\langle\theta\rangle^{2}$ via the uncertainty principle $\Delta\theta\Delta k\geq 1/2$, which yields $\Delta\theta^{2}\geq 1/(N^{2}\mathcal{I})$. Consequently, $\Delta\theta^{2}$ becomes extremely large as $\mathcal{I}\to 0$, while it may be as small as $1/N^{2}$ for $\mathcal{I}\to 1$.
@@
In Fig. 6b the Fisher information is plotted as a function of the boson-boson interaction for $\varepsilon=0$. We notice that at zero temperature, $\mathcal{I}$ is exactly equal to $1/N$ for vanishing boson-boson interaction, while for $U/J$ large and positive it is close to zero. This can be understood from the fact that in such regime the ground state is close to the separable state $|\text{FOCK}\rangle$, so that no large fluctuations of the number of particles in each well are expected. At finite temperature, thermal fluctuations provide a larger probability to other Fock states with a slightly imbalanced population, thus increasing $\mathcal{I}$ (see also Fig. 6a). Conversely, for $U/J$ large and negative $\mathcal{I}$ is close to unity. In this regime the ground state is close to the entangled state $|\text{CAT}_{+}\rangle$, so that in a series of repeated measurements, the observed relative number of particles in the two wells would oscillate violently between zero and $N$. At finite temperature, states with a less prominent particle imbalance acquire a larger probability, resulting in a reduced value of $\mathcal{I}$. In practice, temperature has the tendency of smoothing out the fluctuations of the relative number of particles, so that $\mathcal{I}$ varies more slowly as a function of $U/J$ (Fig. 6b).
@@ III.6 Coherence visibility
$=\frac{1}{Z}\sum_{n,j=0}^{N}e^{-\beta E_{n}}\frac{2}{N}\sqrt{(1+j)(N-j)}c_{j+1}^{(n)}c_{j}^{(n)}$ (29)
@@ Appendix A Validity of the quantum phase model and possible improvements
The number of eigenstates $\psi_{n}(\theta)$ is in principle infinite, reflecting the fact that this approach is naturally suited to describe the system in the thermodynamic limit. Indeed, with canonical quantization the classical observable $k$, bounded in the interval [N/2,+N/2][-N/2,+N/2], is promoted to the unbounded operator $\hat{k}=-i\partial/\partial\theta$, and the mapping between the two can be made rigorous only in the limit $N\to\infty$. However, we can still apply the QPM to describe a system with a finite number $N$ of bosons, by introducing a cutoff on the label $n$ in such a way that the total number of eigenstates (i.e. the dimension of the Hilbert space) is equal to the dimension of the Hilbert space of (II.1), which is $N+1$. If $N$ is even, taking into account the ordering of the eigenvalues in terms of Mathieu characteristic numbers, this means that $n$ must take on the integer values from 0 to $N/2$. Thermal averages may then be computed as
@@
and $Z=\sum_{n=0}^{N/2}e^{-\beta E_{n}^{(a)}}+\sum_{n=1}^{N/2}e^{-\beta E_{n}^{(b)}}$. If instead $N$ is odd, the index $n$ would take on the integer values between 0 and $\lfloor N/2\rfloor$ in the $a$-summations, and the integer values between $1$ and $\lceil N/2\rceil$ in the $b$-summations.